OpenStudy (anonymous):

Pre Calc: Write an equation in standard form for a line with a slope of –1 passing through C(2, 1).

OpenStudy (anonymous):

y-1=(-1)(x-2) y-1= -x+2 y= -x + 3 Point Slope Formula^^^

OpenStudy (anonymous):

@Ren. YOU'RE A LIFE SAVER

OpenStudy (anonymous):

Oh shucks. No problemo though.